# Dimension-independent convergence rates of randomized nets using median-of-means

### v2

@@ Abstract
Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness. Building on this framework, we prove that the median estimator attains dimension-independent convergence under tractability conditions characterized by low effective dimensionality, a property known as strong tractability in complexity theory. Our analysis strengthens existing guarantees by improving the convergence rates and relaxing the theoretical assumptions previously required for dimension-independent convergence.
@@ 1 Introduction
The quasi-Monte Carlo (QMC) method is a powerful alternative to conventional Monte Carlo (MC) techniques for high-dimensional integration. While both methods approximate integrals by averaging $n$ function evaluations, QMC replaces MC’s pseudorandom sampling with low-discrepancy point sets engineered to uniformly explore the sampling space. This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC. These advantages have made QMC particularly valuable in fields like computational finance [13, 18] and physics [16], where high-dimensional integrals are pervasive.
@@
Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets. Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [5]. The random linear scrambling [21], a computationally efficient alternative to Owen’s scrambling [24], preserves the structure of digital nets while introducing controlled randomness. This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference. However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [26].
@@
To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [2, 7, 8, 9, 32]. The median of linearly scrambled digital net estimators was first analyzed in [26], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$. Subsequent work in [27] extended this result to higher dimensions. For finitely differentiable integrands, [25] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
@@
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [17]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [23], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [1], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions. The connection to weighted function spaces [28] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
@@
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [9]. A key distinction lies in the randomization: their framework omits digital shifts (Subsection 2.1), a critical component of our scrambling method, and employs worst-case integration error over the unit ball in a weighted Sobolev space as the primary metric. In contrast, we demonstrate that median estimators incorporating digital shifts achieve near-optimal convergence rates on a fixed integrand with high probability, avoiding worst-case error analysis entirely. This probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [14, 15]. Within this framework, we improve the convergence rates from nearly $O(n^{-\alpha})$ in [9] to nearly $O(n^{-\alpha-1/2})$, where $\alpha$ denotes the smoothness parameter in their analysis (we note that $\alpha$ is defined differently in this paper). Furthermore, we relax the conditions for dimension-independent convergence (Remarks 3 and 9). Specifically, we show digital nets under the completely random design (Subsection 2.1) achieve dimension-independent convergence under conditions comparable to those for polynomial lattices in [9], while digital nets under the random linear scrambling attain the same property under strictly weaker conditions than those assumed in [9].
@@ 2 Background and notation
Let =$\mathbb{N}=\{1,2,3,\dots\}$ denote the natural numbers and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $s$-dimensional indices, we define $\mathbb{N}^{s}_{*}=\mathbb{N}_{0}^{s}\setminus\{\boldsymbol{0}\}$ to exclude the zero vector. We define =$\mathbb{Z}_{\leqslant\ell}=\{0,1,\dots,\ell\}$ for $\ell\in\mathbb{N}_{0}$ and <=$\mathbb{Z}_{<\ell}=\{0,1,\dots,\ell-1\}$ for $\ell\in\mathbb{N}$. The dimension of the integration domain is denoted by $s$, and we write $1{:}s=\{1,2,\dots,s\}$ for $s\in\mathbb{N}$. For a vector $\boldsymbol{x}\in{}^{s}$ and a subset $u\subseteq 1{:}s$, $\boldsymbol{x}_{u}$ denotes the subvector of $\boldsymbol{x}$ indexed by $u$, while $\boldsymbol{x}_{u^{c}}$ denotes the subvector indexed by $1{:}s\setminus u$. The indicator function $\boldsymbol{1}\{\mathcal{A}\}$ equals $1$ if event $\mathcal{A}$ occurs and 00 otherwise. For $\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{s})\in\mathbb{N}^{s}_{0}$, $\|\boldsymbol{\ell}\|_{1}=\sum_{j=1}^{s}\ell_{j}$. The cardinality of a set $K$ is $|K|$. For positive sequence $a_{n}$ and $b_{n}$, $a_{n}=O(b_{n})$ denotes asymptotic dominance: $a_{n}\leqslant Cb_{n}$ for some constant $C<\infty$ and all sufficiently large $n\in\mathbb{N}$.
@@ 2.1 Digital net construction and randomization
For $m\in\mathbb{N}$ and $i\in\mathbb{Z}_{2^{m}}$, we express $i$ in its binary form $i=\sum_{\ell=1}^{m}i_{\ell}2^{\ell-1}$, encoded as the column vector $\vec{i}=(i_{1},\dots,i_{m})^{\mathsf{T}}\in\{0,1\}^{m}$. Analogously, for $a\in[0,1)$, we approximate its infinite binary expansion $a=\sum_{\ell=1}^{\infty}a_{\ell}2^{-\ell}$ by truncating to $E\in\mathbb{N}$ digits, yielding $\vec{a}=(a_{1},\dots,a_{E})^{\mathsf{T}}\in\{0,1\}^{E}$. For numbers with dual representations (such as dyadic rationals), we select the finite expansion ending in zeros.
A base-2 digital net in $[0,1]^{s}$ is determined by $s$ matrices $C_{j}\in\{0,1\}^{E\times m}$. The deterministic points $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{is})$ are constructed via:
$\vec{x}_{ij}=C_{j}\vec{i}\ \ \mathsf{mod}\ 2\text{ for }i\in\mathbb{Z}_{2^{m}},j\in 1{:}s,$
@@
[…]
where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$. The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli$(0.5)$ entries. A common way to randomize $C_{j}$ is the random linear scrambling [21]:
[…]
where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli$(0.5)$), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [5, Chapter 4.4] for details).
While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis. We address this in Section 5. For initial insights, Section 4 adopts the complete random design [25], where all entries of $C_{j}$ are drawn independently from Bernoulli$(0.5)$. This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
We conveniently assume $E=\infty$ and focus our theoretical analysis on $\hat{\mu}_{\infty}$. In practice, $E$ is constrained by computational precision. Lemma 1 of [27] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$. Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
[…]
@@
The equality holds in the $L^{2}$-sense. Using this decomposition, [25, Theorem 1] establishes a key error structure for QMC estimators:
@@
2.3 Probabilistic error guarantee of the median
The analysis of equation (4) uses two key properties of the coefficients $S(\boldsymbol{k})$: they are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [26, Lemma 4] for the univariate case $s=1$). These properties allow us to compute the variance of $\hat{\mu}_{\infty}$ and construct probabilistic error bounds by Chebyshev’s inequality.
To sharpen these bounds, we introduce a low-probability exclusion event $\mathcal{A}$ satisfying $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and $\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\ll\mathrm{Var}(\hat{\mu}_{\infty})$. This leads to the following median concentration result:
@@ 2.3 Probabilistic error guarantee of the median
Let $\mathcal{A}$ be an event with $\Pr(\mathcal{A})\leqslant\delta<1/8$ and $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$. Then the sample median $\hat{\mu}^{(r)}_{\infty}$ of $12r-$ 1 independently generated replicates of $\hat{\mu}_{\infty}$ satisfies
$\Pr\Big{(}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\Big{)}\leqslant(8\delta)^{r}.$
@@
This is an intermediate result proven in [27, Lemma 6]. ∎
For $f\in C([0,1]^{s})$, the above lemma implies the mean squared error bound
$\mathbb{E}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}\leqslant\delta^{-1}\Pr(\mathcal{A}^{c})\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})+(8\delta)^{r}\|f\|^{2}_{\infty}.$
@@
We choose $\mathcal{A}$ to be the event $\{Z(\boldsymbol{k})=1$ for some $\boldsymbol{k}\in K\}$ for a carefully chosen index set $K\subseteq\mathbb{N}^{s}_{*}$. It is shown in [25, Subsection 2.4] that
${\Pr(\mathcal{A}^{c})}\mathrm{Var}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})\leqslant\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$ (5)
The central challenge is to select $K$ such that this sum converges at the desired rate. Our analysis addresses this by constructing $K$ adaptively based on the decay rate of the Walsh coefficients $\hat{f}(\boldsymbol{k})$, which encodes the integrand’s smoothness and effective dimensionality.

### v3

@@ Abstract
Recent advances in quasi-Monte Carlo integration demonstrate that the median of linearly scrambled digital net estimators achieves near-optimal convergence rates for high-dimensional integrals without requiring a priori knowledge of the integrand’s smoothness. Building on this framework, we prove that the median estimator attains dimension-independent convergence, a property known as strong tractability in complexity theory, under tractability conditions characterized by low effective dimensionality. Using a probabilistic, integrand-specific error criterion, our analysis establishes both faster and dimension-independent convergence under weaker assumptions than previously possible in the worst-case setting.
@@ 1 Introduction
The quasi-Monte Carlo (QMC) method is a powerful alternative to conventional Monte Carlo (MC) techniques for high-dimensional integration. While both methods approximate integrals by averaging $n$ function evaluations, QMC replaces MC’s pseudorandom sampling with low-discrepancy point sets engineered to uniformly explore the sampling space. This structured approach allows QMC to avoid the exponential cost growth with dimension of classical quadrature rules while achieving faster error convergence than MC. These advantages have made QMC particularly valuable in fields like computational finance [joy1996quasi, l2009quasi] and physics [kuo:nuye:2016], where high-dimensional integrals are pervasive.
@@
Building on these strengths, we focus on randomized quasi-Monte Carlo (RQMC) methods based on linearly scrambled base-2 digital nets. Digital nets are a widely studied QMC construction that achieves low-discrepancy sampling through deterministic point sets [dick:pill:2010]. The random linear scrambling [mato:1998:2], a computationally efficient alternative to Owen’s scrambling [rtms], preserves the structure of digital nets while introducing controlled randomness. This randomization enhances the root mean squared error convergence rate to nearly $O(n^{-3/2})$ under smoothness conditions and enables statistical inference. However, traditional approaches that estimate integrals by averaging independent QMC replicates face a critical limitation: outliers inherent to linearly scrambled digital nets can prevent these methods from attaining faster convergence rates, even when the integrand satisfies stronger smoothness assumptions [superpolyone].
@@
To address this limitation, we adopt the median trick, a methodological innovation that has inspired many recent advances in high-dimensional integration [chen2025randomintegrationalgorithmhighdimensional, goda2024simpleuniversalalgorithmhighdimensional, goda:lecu:2022, Goda2024, ye2025medianqmcmethodunbounded]. The median of linearly scrambled digital net estimators was first analyzed in [superpolyone], which demonstrated a convergence rate of $O(n^{-c\log(n)})$ for one-dimensional analytic functions on $[0,1]$, where $c<3\log(2)/\pi^{2}\approx 0.21$. Subsequent work in [superpolymulti] extended this result to higher dimensions. For finitely differentiable integrands, [pan2024automatic] further showed that the median estimator achieves near-optimal asymptotic convergence rates across diverse function spaces, requiring neither a priori smoothness parameters nor pre-optimized digital nets.
@@
A critical gap in these prior studies is the role of dimensionality. We focus on the regime where the dimension grows unboundedly while the function norm remains bounded—a common scenario in path integral approximations [kuo:sloa:wasi:wozn:2010]. Our analysis demonstrates that the median estimator attains dimension-independent convergence, a property termed strong tractability in information-based complexity [nova:wozn:2008], under assumptions on the decay of the relative variation $\gamma_{u}$ of each ANOVA component (formally defined in Section 3). This notion of relative variation aligns with the effective dimension framework introduced in [cafmowen], which quantifies how high-dimensional integrands can concentrate their variation in low-order interactions. The connection to weighted function spaces [sloa:wozn:1998] is discussed in Remark 1, further bridging our tractability conditions to classical QMC theory.
@@
Before presenting our results, we compare our contributions to the tractability analysis of median digital nets in [Goda2024]. A key distinction lies in the error criterion: their analysis employs the worst-case integration error over the unit ball in a weighted Sobolev space, whereas we demonstrate that median estimators achieve near-optimal convergence rates on a fixed integrand with high probability. Our probabilistic approach is justified by the exponential decay in the likelihood of large errors as the number of replicates increases, aligning with the $(\varepsilon,\delta)$-approximation framework studied in [kunsch2019solvable, kunsch2019optimal].
Within this framework, we prove, informally stated, the following result:
Theorem 1.
Given $\eta>0$, $k\in\mathbb{N}$ and $f\in C^{(k,\dots,k)}([0,1]^{s})$, it holds with probability $1-2^{-r}$ that the median estimator $\hat{\mu}^{(r)}_{\infty}$ from Subsection 2.3 satisfies
$|\hat{\mu}^{(r)}_{\infty}-\mu|\leqslant C_{\eta,k,f}n^{-k-1/2+\eta}.$
The constant $C_{\eta,k,f}$ is independent of $s$ under the random linear scrambling if the relative variation $\gamma_{u}$ satisfies the tractability conditions in Theorem 5.
These nearly $O(n^{-k-1/2})$ integrand-specific error rates complement the nearly $O(n^{-k})$ worst-case rates established in [Goda2024]. Furthermore, Theorem 5 requires strictly weaker assumptions than those assumed in [Goda2024] for the random linear scrambling, demonstrating that dimension-independent convergence is more easily achieved under our error criterion.
As a final remark, we note that while several algorithms match lower complexity bounds for integration in specific spaces—such as higher-order digital nets [goda2016explicit, goda2017optimal, goda2018optimal] and Frolov lattice rules [frolov1976upper, krieg2017universal, ullrich2016role]—our method offers distinct practical advantages. It requires no smoothness parameter specification, yet automatically matches theoretical lower bounds up to an arbitrarily small $\eta>0$ in the convergence rate, making it both efficient and immune to parameter misspecification. Combined with its straightforward implementation and the dimension-independent convergence established in this paper, the median estimator presents a highly practical and robust alternative for applied problems. Furthermore, recent work has successfully extended median RQMC to $L^{2}$-approximation [pan2025universal, pan2025l_2], demonstrating its broader utility beyond numerical integration.
@@ 2 Background and notation
Let $\mathbb{N}=\{1,2,3,\dots\}$ denote the natural numbers and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $s$-dimensional indices, we define $\mathbb{N}^{s}_{*}=\mathbb{N}_{0}^{s}\setminus\{\boldsymbol{0}\}$ to exclude the zero vector. We define $\mathbb{Z}_{\leqslant\ell}=\{0,1,\dots,\ell\}$ for $\ell\in\mathbb{N}_{0}$ and $\mathbb{Z}_{<\ell}=\{0,1,\dots,\ell-1\}$ for $\ell\in\mathbb{N}$. The dimension of the integration domain is denoted by $s$, and we write $1{:}s=\{1,2,\dots,s\}$ for $s\in\mathbb{N}$. For a vector $\boldsymbol{x}\in\real^{s}$ and a subset $u\subseteq 1{:}s$, $\boldsymbol{x}_{u}$ denotes the subvector of $\boldsymbol{x}$ indexed by $u$, while $\boldsymbol{x}_{u^{c}}$ denotes the subvector indexed by $1{:}s\setminus u$. The indicator function $\boldsymbol{1}\{\mathcal{A}\}$ equals $1$ if event $\mathcal{A}$ occurs and 0 otherwise. For $\boldsymbol{\ell}=(\ell_{1},\dots,\ell_{s})\in\mathbb{N}^{s}_{0}$, $\|\boldsymbol{\ell}\|_{1}=\sum_{j=1}^{s}\ell_{j}$. The cardinality of a set $K$ is $|K|$. For positive sequences $a_{n}$ and $b_{n}$, $a_{n}=O(b_{n})$ denotes asymptotic dominance: $a_{n}\leqslant Cb_{n}$ for some constant $C<\infty$ and all sufficiently large $n\in\mathbb{N}$.
@@ 2.1 Digital net construction and randomization
For $m\in\mathbb{N}$ and $i\in\mathbb{Z}_{<2^{m}}$, we express $i$ in its binary form $i=\sum_{\ell=1}^{m}i_{\ell}2^{\ell-1}$, encoded as the column vector $\vec{i}=(i_{1},\dots,i_{m})^{\mathsf{T}}\in\{0,1\}^{m}$. Analogously, for $a\in[0,1)$, we approximate its infinite binary expansion $a=\sum_{\ell=1}^{\infty}a_{\ell}2^{-\ell}$ by truncating to $E\in\mathbb{N}$ digits, yielding $\vec{a}=(a_{1},\dots,a_{E})^{\mathsf{T}}\in\{0,1\}^{E}$. For numbers with dual representations (such as dyadic rationals), we select the finite expansion ending in zeros.
A base-2 digital net in $0,1]^{s}$ is determined by $s$ matrices $C_{j}\in\{0,1\}^{E\times m}$. The deterministic points $\boldsymbol{x}_{i}=(x_{i1},\dots,x_{is})$ are constructed via:
$\vec{x}_{ij}=C_{j}\vec{i}\ \ \mathsf{mod}\ 2\text{ for }i\in\mathbb{Z}_{<2^{m}},j\in 1{:}s,$
@@
$\vec{x}_{ij}=C_{j}\vec{i}+\vec{D}_{j}\ \ \mathsf{mod}\ 2,$ (1)
where $C_{j}\in\{0,1\}^{E\times m}$ and $\vec{D}_{j}\in\{0,1\}^{E}$ are random with precision $E\geqslant m$. The vector $\vec{D}_{j}$ is called the digital shift and comprises independent Bernoulli(0.5)(0.5) entries. A common way to randomize $C_{j}$ is the random linear scrambling [mato:1998:2]:
$C_{j}=M_{j}\mathcal{C}_{j}\ \ \mathsf{mod}\ 2,$
where $M_{j}\in\{0,1\}^{E\times m}$ is a random lower-triangular matrix (diagonal entries fixed at $1$, subdiagonal entries independently drawn from Bernoulli(0.5)(0.5)), and $\mathcal{C}_{j}\in\{0,1\}^{m\times m}$ is a predetermined generating matrix optimized to improve the discrepancy of resulting points $\boldsymbol{x}_{i}$ (see [dick:pill:2010, Chapter 4.4] for details).
While effective, the reliance of random linear scrambling on carefully designed generating matrices complicates theoretical analysis. We address this in Section 5. For initial insights, Section 4 adopts the complete random design [pan2024automatic], where all entries of $C_{j}$ are drawn independently from Bernoulli(0.5)(0.5). This approach retains the asymptotic convergence rate of random linear scrambling while circumventing the need for pre-optimized generator matrices.
@@
We conveniently assume $E=\infty$ and focus our theoretical analysis on $\hat{\mu}_{\infty}$. In practice, $E$ is constrained by computational precision. Lemma 1 of [superpolymulti] shows $|\hat{\mu}_{E}-\hat{\mu}_{\infty}|\leqslant\omega_{f}(\sqrt{s}2^{-E})$, where $\omega_{f}$ is the modulus of continuity of $f$. Thus, $\hat{\mu}_{E}$ closely approximates $\hat{\mu}_{\infty}$ when $\omega_{f}(\sqrt{s}2^{-E})\ll|\hat{\mu}_{E}-\mu|$.
@@ 2.2 Walsh functions and error analysis for base-2 digital nets
where $\boldsymbol{k}=(k_{1},\dots,k_{s})\in\mathbb{N}^{s}_{0}$. These functions constitute a complete orthonormal basis for $L^{2}([0,1]^{s})$ [dick:pill:2010], enabling the Walsh series decomposition:
@@
The equality holds in the $L^{2}$-sense. Using this decomposition, [pan2024automatic, Theorem 1] establishes a key error structure for QMC estimators:
@@
2.3 Probabilistic error guarantee of the median estimator
Given $r\in\mathbb{N}$, we consider 2r12r-1 independently generated replicates of $\hat{\mu}_{\infty}$ ordered from the smallest to the largest:
$\hat{\mu}^{(1)}_{\infty}\leqslant\dots\leqslant\hat{\mu}^{(r)}_{\infty}\leqslant\dots\leqslant\hat{\mu}^{(2r-1)}_{\infty}.$
A key property of the sample median $\hat{\mu}^{(r)}_{\infty}$ is the following probabilistic error bound:
@@ 2.3 Probabilistic error guarantee of the median estimator
Let $K\subseteq\mathbb{N}^{s}_{*}$ and $\mathcal{A}$ be the event $\{Z(\boldsymbol{k})=1\text{ for some }\boldsymbol{k}\in K\}$. If $\Pr(\mathcal{A})\leqslant\delta<1/8$, then
$\Pr(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant(8\delta)^{r}\quad\text{for}\quad\varepsilon_{K}=\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1)|\hat{f}(\boldsymbol{k})|^{2}.$
@@
First note that the random signs $S(\boldsymbol{k})$ in equation (4) are zero-mean and pairwise-independent for distinct $\boldsymbol{k}\in\mathbb{N}_{*}^{s}$ (see [superpolyone, Lemma 4] for the one-dimensional case). Since $\mathcal{A}$ is independent of $S(\boldsymbol{k})$, we have $\mathbb{E}(\hat{\mu}_{\infty}\!\mid\!\mathcal{A}^{c})=\mu$ and
$\mathrm{Var}(\hat{\mu}_{\infty}\mid\mathcal{A}^{c})=\sum_{\boldsymbol{k}\in\mathbb{N}^{s}_{*}\setminus K}\Pr(Z(\boldsymbol{k})=1\mid\mathcal{A}^{c})|\hat{f}(\boldsymbol{k})|^{2}\leqslant\frac{\varepsilon_{K}}{\Pr(\mathcal{A}^{c})}.$
Applying Chebyshev’s inequality yields
$\Pr(|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant$ $\Pr(\mathcal{A})+\Pr(\mathcal{A}^{c})\Pr(|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta\mid\mathcal{A}^{c})\leqslant 2\delta.$
Finally, $|\hat{\mu}^{(r)}_{\infty}-\mu|>\varepsilon_{K}/\delta$ only if the event $|\hat{\mu}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta$ occurs for at least $r$ out of the 2r12r-1 independent replicates. A union bound then implies
$\Pr(|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}>\varepsilon_{K}/\delta)\leqslant{2r-1\choose r}(2\delta)^{r}\leqslant 2^{2r-1}(2\delta)^{r}\leqslant(8\delta)^{r}.\qed$
This lemma justifies the use of the sample median $\hat{\mu}^{(r)}_{\infty}$ as an estimator for $\mu$. For $f\in C([0,1]^{s})$, the mean squared error satisfies
$\mathbb{E}|\hat{\mu}^{(r)}_{\infty}-\mu|^{2}\leqslant\varepsilon_{K}/\delta+(8\delta)^{r}\|f\|^{2}_{\infty}.$
@@
The main challenge is therefore to select a set $K$ so that $\varepsilon_{K}$ converges at the desired rate. We address this by constructing $K$ adaptively based on the decay rate of the Walsh coefficients $\hat{f}(\boldsymbol{k})$, which encodes the integrand’s smoothness and effective dimensionality. Notably, the median estimator $\hat{\mu}^{(r)}_{\infty}$ does not require explicit knowledge of $K$, making it applicable without prior information.